OpenStudy (anonymous):

Solve the equation 5^(x-1)=3^x round answer to 3 decimal places

OpenStudy (dumbcow):

take log both sides \[\ln 5^{x-1} = \ln 3^{x}\] \[(x-1) \ln 5 = x \ln 3\] \[x \ln 5 - \ln 5 = x \ln 3\] \[x \ln 5 - x \ln 3 = \ln 5\] \[x = \frac{\ln 5}{\ln 5 - \ln 3}\]

OpenStudy (anonymous):

i got 3.151

OpenStudy (dumbcow):

yep